Experiment 26: The Solubility Product of Ba(IO3)2 : Processing the DataExperiment 26 Data and Calculations: The Solubility Product of Ba(IO3)2 The calculations here are similar to those in Experim218 Experiment 26 Determination of the Solubility Product of Ba(10) Test Tube Number Moles Ba? precipitated Moles Bain equil,
Answer #1

Initial moles Ba2+

Moles Ba2+ = (volume of Ba(NO3)2 in L)*(molarity of Ba(NO3)2)

Take test tube 1 as an example.

Moles Ba2+ = (5.00 mL)*(1 L)/(1000 mL)*(0.0200 M)

= 1.00*10-4 mol

Fill in the table.

Test Tube Number

1

2

3

4

5

Initial mols Ba2+

1.00*10-4

1.00*10-4

1.00*10-4

1.00*10-4

1.00*10-4

Initial moles IO3-

Moles IO3- = (volume of KIO3 in L)*(molarity of KIO3)

Take test tube 1 as an example.

Moles IO3- = (1.00 mL)*(1 L)/(1000 mL)*(0.0350 M)

= 3.50*10-5 mol.

Fill in the table.

Test Tube Number

1

2

3

4

5

Initial mols IO3-

3.50*10-5

7.00*10-5

1.05*10-4

1.40*10-4

1.75*10-4

Moles IO3- in solution

Moles IO3- = (volume of solution in L)*(calculated molarity of IO3- in solution)

Take test tube 1 as an example.

Molarity IO3- in solution = 1.025*10-3 M.

Moles IO3- in solution in test tube 1 = (12.00 mL)*(1 L)/(1000 mL)*(1.025*10-3 M)

= 1.23*10-5 mol

Fill in the table.

Test Tube Number

1

2

3

4

5

Mols IO3- in solution

1.23*10-5

1.2756*10-5

≈ 1.27*10-5

1.3008*10-5

≈ 1.30*10-5

1.5912*10-5

≈ 1.59*10-5

1.9836*10-5

≈ 1.98*10-5

Moles IO3- precipitated

Moles IO3- precipitated = (initial mols IO3-) – (mols IO3- in solution)

Take test tube 1 as an example.

Moles IO3- precipitated = (3.50*10-5) – (1.23*10-5)

= 2.27*10-5

Fill in the table.

Test Tube Number

1

2

3

4

5

Mols IO3- precipitated

2.27*10-5

5.73*10-5

9.20*10-5

1.241*10-4

≈ 1.24*10-4

1.552*10-4

≈ 1.55*10-4

Moles Ba2+ precipitated

Consider the ionization of Ba(IO3)2.

Ba(IO3)2 (s) <=======> Ba2+ (aq) + 2 IO3- (aq)

As per the stoichiometric equation above,

1 mol Ba2+ = 2 mols IO3-

Therefore,

Mols Ba2+ precipitated = ½*(mols IO3- precipitated)

Take test tube 1 as an example.

Mols Ba2+ precipitated = ½*(2.27*10-5 mol)

= 1.135*10-5

Fill in the table.

Test Tube Number

1

2

3

4

5

Mols Ba2+ precipitated

1.135*10-5

2.865*10-5

4.60*10-5

6.20*10-5

7.75*10-5

Moles Ba2+ in equilibrium solution

Moles Ba2+ in equilibrium solution = (initial mols Ba2+) – (mols Ba2+ precipitated)

Take test tube as an example.

Mols Ba2+ in equilibrium solution = (1.00*10-4) – (1.135*10-5)

= 8.865*10-5

Fill in the table.

Test Tube Number

1

2

3

4

5

Mols Ba2+ in equilibrium solution

8.865*10-5

7.135*10-5

5.40*10-5

3.80*10-5

2.25*10-5

[Ba2+] in equilibrium solution

Volume of the equilibrium solution = 12.00 mL = (12.00 mL)*(1 L)/(1000 mL)

= 0.012 L

[Ba2+] = (mols Ba2+ in the equilibrium solution)/(0.012 L)

= 7.3875*10-3 M

≈ 7.39*10-3 M

Fill in the table below.

Test Tube Number

1

2

3

4

5

[Ba2+] in equilibrium solution, M

7.39*10-3

5.94*10-3

4.50*10-3

3.16*10-3

1.875*10-3

≈ 1.87*10-3

Ksp for Ba(IO3)2

Ksp = [Ba2+][IO3-]2

Take test tube 1 as an example.

Ksp = (7.39*10-3)*(1.025*10-3)2

= 7.76*10-9 (ignore units)

Fill in the table below.

Test Tube Number

1

2

3

4

5

Ksp

7.76*10-9

6.71*10-9

5.29*10-9

5.56*10-9

5.11*10-9

Mean value of Ksp

Mean value = 1/5*(7.76 + 6.71 + 5.29 + 5.56 + 5.11)*10-9

= 6.086*10-9

≈ 6.09*10-9

Standard deviation of Ksp = √Σ(x - x̅)2/(n – 1)

where x̅ is the mean of the Ksp values and n is the number of measurements.

Standard Ksp = 1.12*10-9 (ignore units)